Ardoin Enterprises has to determine its cost of capital using the following information:

The firm has $40,000,000 in corporate bonds currently selling at 97.5. The bonds mature in 9 years and have an annual coupon rate of 6.6% paid semiannually. The firm faces a 34% tax rate and has 1,500,000 shares of preferred stock that pays a dividend of $0.80 per year and currently sells for $9.00 per share.

Common stock selling for $3.50 per share has just paid a dividend of $0.30 and is expected to grow by 4% forever. The firm has a beta of 1.4 and the risk free rate on treasury securities is 2.5%. The average return on the S&P500 is 12.54%.

Calculate the cost of capital for the firm. Ardoin Enterprises has 25 million common shares outstanding.
Answer #1
As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 2.5 + 1.4 * (12.54 - 2.5)
Expected return% = 16.56
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
3.5 = 0.3 * (1+0.04) / (Cost of equity - 0.04)
Cost of equity% = 12.91

Avg cost of equity = (16.56+12.91)/2 = 14.74%

MV of equity=Price of equity*number of shares outstanding
MV of equity=3.5*25000000
=87500000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*40000*0.975
=39000000
MV of firm = MV of Equity + MV of Bond
=87500000+39000000
=126500000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 87500000/126500000
W(E)=0.6917
Weight of debt = MV of Bond/MV of firm
Weight of debt = 39000000/126500000
W(D)=0.3083
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =9x2
975 =∑ [(6.6*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^9x2
                   k=1
YTM = 6.9787503947
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 6.9787503947*(1-0.34)
= 4.605975260502
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=4.61*0.3083+14.74*0.6917
WACC =11.62%